Twice of a number is greater than half of the same number by 60.Find the number.
1
Expert's answer
2012-01-05T08:19:07-0500
Let's denominate the unknown number by N. Now we can write down an equation:
2N - 60 = N/2.
So, 1.5N = 60 ==> N = 60/1.5 = 40.